A music store is selling compact discs for $11.50 and $7.50. You buy 12 discs and spend a total of $106.00. How many compact dics that cost 11.50 did you buy. Show all work and write an equation
Answer by checkley75(3666)

11.50x+7.50(12-x)=106
11.50x+90-7.5x=106
4x=106-90
4x=16
x=16/4
x=4 ans. for the number of compact discs bought.
Proof:
11.50*4+7.5(12-4)=106
46+7.5*8=106
46+60=106
106=106